### v1

# multiGradICON: A Foundation Model for Multimodal Medical Image Registration

Abstract
Modern medical image registration approaches predict deformations using deep networks. These approaches achieve state-of-the-art (SOTA) registration accuracy and are generally fast. However, deep learning (DL) approaches are, in contrast to conventional non-deep-learning-based approaches, anatomy-specific. Recently, a universal deep registration approach, uniGradICON, has been proposed. However, uniGradICON focuses on monomodal image registration. In this work, we therefore develop multiGradICON as a first step towards universal multimodal medical image registration. Specifically, we show that 1) we can train a DL registration model that is suitable for monomodal and multimodal registration; 2) loss function randomization can increase multimodal registration accuracy; and 3) training a model with multimodal data helps multimodal generalization. Our code and the multiGradICON model are available at https://github.com/uncbiag/uniGradICON.
Keywords:
Medical image registration Deep learning Multimodal.
1 Introduction
Learning-based medical image registration [9, 53, 59, 5] has significantly improved over the last decade. Current SOTA learning-based deep registration networks [48, 47, 35, 36] are faster and more accurate than conventional registration methods using numerical optimization. However, learning-based approaches generally 1) focus on monomodal image registration and 2) are trained for a specific anatomical region (often the brain), making them much less generically applicable than conventional approaches. While many monomodal learning-based approaches have been developed, uniGradICON [47] is currently the only learning-based approach that is designed to be a universal method supporting registrations for different anatomies in one model working directly with images and SynthMorph [23] may generalize to other anatomies by training on synthetic shapes. Further, uniGradICON focuses on monomodal registration, and multimodal generalization is primarily achieved by instance optimization (IO). Hence, our goal is to further close the gap between conventional and learning-based registration approaches by developing multiGradICON, a multimodal generalization of uniGradICON. multiGradICON extends uniGradICON by 1) choosing a multimodal similarity measure, 2) incorporating multimodal registration tasks into training, and 3) exploring monomodal, multimodal, and randomized strategies for image similarity loss based on multimodal registration network inputs.
Our specific contributions are to show that
  • 1)
    while uniGradICON is a strong baseline for monomodal registration, even for unseen modalities, it does not generalize well to multimodal registration when modalities are drastically different (see Fig. 1 for an example);
  • 2)
    training a monomodal model with a squared local normalized cross correlation image similarity loss ($1-\text{LNCC}^{2}$) does not lead to multimodal generalization in the absence of multimodal registration tasks during training;
  • 3)
    multiGradICON allows for multimodal generalization while retaining good monomodal registration accuracy;
  • 4)
    similarity loss randomization (i.e., randomly picking which modalities should be compared in the loss) improves registration accuracy for datasets containing multi-parametric data, even when scalar images are used for inference.
Figure 1: Comparison of uni- and multiGradICON on T1w MRI-mean diffusivity (MD) registration from ABCD. Note the much better matching of the ventricles for multiGradICON.
2 Related work
Following uniGradICON [47] (which builds on ICON [15] and GradICON [48]), we focus on non-parametric image registration [34] where a displacement field is predicted. Such image registrations are generally formulated as a balance between an image (dis)similarity term and a regularizer encouraging spatial smoothness [34]. By extending uniGradICON, we retain its gradient inverse consistent regularizer (see Sec. 3) and focus on multimodal alternatives to the negative local normalized cross correlation ($1-\text{LNCC}$) loss used in uniGradICON.
Multimodal similarity measures. Multimodal registration requires similarity measures that extend beyond the direct intensity similarities measured by mean squared or absolute error (MSE/MAE) losses. The goal is to assess whether an image pair is spatially aligned even though image appearance may be quite different across modalities. Conventional approaches maximize measures of statistical dependence such as normalized cross correlation (NCC) or squared NCC (if the sign of the correlation is unknown). Even more general statistical dependencies can be captured by maximizing mutual information (MI) between image pairs [51]. Allowing for more local control, these measures have been extended to local NCC (LNCC; as in uniGradICON) or its squared variant as well as local MI; see [22, 34] for an overview of these conventional similarity measures. Such conventional similarity measures have been used to train multimodal deep registration networks [48, 42, 16]. More modern approaches target similarity measures depending on image self-similarity (such as the modality independent neighborhood descriptor (MIND) [20] and its improved version with self-similarity context (MIND-SSC) [21]) or local entropy images [52]. The general idea of these more modern similarity measures is to sidestep image differences by using a representation that remains similar even if the underlying image pairs look different. On the extreme end of the similarity measure spectrum, one can then use image segmentations via a Dice loss or try to transform one modality into the other via image synthesis. As such segmentations are generally not directly available and the intensity relationship between image pairs is not a-priori known learning-based approaches are commonly employed.
Learning-based approaches for multimodal similarity quantification. Learning better multimodal similarity measures has been explored via non-deep-learning [30] and deep learning approaches [44, 57, 10, 43, 49, 31, 37]. These multimodal similarity measures can then be used for conventional optimization-based image registration or to train a multimodal deep registration network. In general, multimodal registration can be simplified by 1) converting the multimodal problem into a monomodal one via image synthesis [54, 41, 39]; or 2) if already aligned images of the same subject are available (e.g., multi-parametric sequences in magnetic resonance imaging (MRI)) by only using monomodal pairings for the similarity loss between subjects but multimodal pairings as the input to the learning formulation [58, 6]; or by 3) training a segmentation model on both modalities and then using the segmentations for the similarity loss [13]. Several works use the Dice loss computed from image segmentations to train a multimodal registration network [26, 45, 29, 23]. Here, segmentations are either obtained manually [26, 45], via automatic segmentation algorithms [29], or via label-image pair synthesis [23]. See the recent review articles on deep-learning-based image registration for further details on multimodal formulations [9, 53].
multiGradICON can use any multimodal similarity measure. However, as a first step, we focus on $1-\text{LNCC}^{2}$ as the similarity training loss as it is closely related to uniGradICON’s $1-\text{LNCC}$ loss. We use $1-\text{LNCC}^{2}$ as well as MIND-SSC [21] for instance optimization. This allows us to keep experiments simple.
3 Methodology
We follow the uniGradICON approach with three key differences: 1) we adjust the image similarity measure so that it is appropriate for multimodal image registration, 2) we use a larger training dataset that contains multimodal registration tasks, and 3) we explore image similarity loss randomization. We introduce the uniGradICON methodology below and highlight our modifications.
Notation. We denote our source and target images by $\left(I^{A},I^{B}\right)$, and consider them to be functions from an image domain to the real numbers. We denote a registration network by $\Phi_{\theta}$. This network $\Phi_{\theta}$ operates on a pair of images to yield a function $\Phi_{\theta}[I^{A},I^{B}]:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which, when precomposed with the source image, is intended to align the images: $I^{A}\circ\Phi_{\theta}[I^{A},I^{B}]\sim I^{B}\,.$
3.1 Architecture
[…]
$\texttt{TS}\{\Psi_{\theta}^{1},\Psi_{\theta}^{2}\}[I^{A},I^{B}]$ $:=\Psi_{\theta}^{1}[I^{A},I^{B}]\circ\Psi_{\theta}^{2}[I^{A}\circ\Psi_{\theta}^{1}[I^{A},I^{B}],I^{B}]\,,$ (1)
$\texttt{DS}\{\Psi_{\theta}\}[I^{A},I^{B}]$ $:=\Psi_{\theta}[\texttt{averagePool}(I^{A},2),\texttt{averagePool}(I^{B},2)]\,.$ (2)
Specifically, we use U-Nets [11] ($\Psi_{\theta}^{i}$) that predict displacement fields and construct the registration network as $\Phi_{\theta}:=\texttt{TS}\{\texttt{TS}\{\texttt{DS}\{\texttt{TS}\{\texttt{DS}\{\Psi_{\theta}^{1}\},\Psi_{\theta}^{2}\}\},\Psi_{\theta}^{3}\},\Psi_{\theta}^{4}\}\,.$
3.2 Training losses
Baseline. The training loss proposed in GradICON [48] is
$\mathcal{L}{=}\mathcal{L}_{\operatorname{sim}}\left(I^{A}\circ\Phi_{\theta}\left[I^{A},I^{B}\right],I^{B}\right)$ $+\mathcal{L}_{\operatorname{sim}}\left(I^{B}\circ\Phi_{\theta}\left[I^{B},I^{A}\right],I^{A}\right)$ (3)
$+\lambda\left\|\nabla\left(\Phi_{\theta}\left[I^{A},I^{B}\right]\circ\Phi_{\theta}\left[I^{B},I^{A}\right]\right)-\mathbf{I}\right\|_{F}^{2}.$
We use the negative squared localized normalized cross correlation ($1-\text{LNCC}^{2}$) as image similarity measure, $\mathcal{L}_{\text{sim}}(\cdot,\cdot)$, in contrast to uniGradICON’s ($1-\text{LNCC}$) loss which assumes locally positive correlations. In contrast, ($1-\text{LNCC}^{2}$) is agnostic to the signs of the correlations. Therefore, it is more appropriate for general multimodal image registration where it is unclear if image intensity pairs are positively or negatively correlated locally.
Image similarity loss randomization. Some multimodal medical datasets provide different modalities for the same patient and anatomical region, e.g., for multi-parametric MRI where images from multiple MR sequences are available (this is, for example, the case for brain MRIs of the Human Connectome Project (HCP) or from BratsReg; see Tab. 1 for details) [^1] For notational simplicity, we denote multiple MR sequences also as multimodal rather than multi-sequence data.. These within-patient images are already aligned since they are derived from the same acquisition. We propose a training strategy to further benefit from these paired multimodal images.
Most deep registration approaches first predict the transformation map to warp the source image to the space of the target image. Image similarity is then calculated between the warped source image and the target image. If the source and target image come from a different modality, this will require a multimodal image similarity measure. If each patient has multiple paired images of different modalities, then the two simplest possible extensions of this approach are to 1) pick one specific modality per patient and proceed with a multimodal image similarity loss or 2) if all patients have the same set of modalities to simply use vector-valued images. The latter approach complicates training a universal model as it would require all datasets to share the same set of modalities which is unrealistic. Instead, we propose extending the former approach. However, we do not pick a specific modality per patient but rather pick a random modality per patient as the input to the deep registration network and another random one to compute the multimodal similarity loss. In expectation, this strategy will train the network with all possible input and loss combinations.
Formally, we assume that our dataset $D=\{P_{i}=\{I_{i}^{1},I_{i}^{2},...,I_{i}^{m}\}\}_{i\in[n]}$ consists of $m$ scans in different modalities for each patient $P_{i}$. We first sample a patient pair $\left(P_{A},P_{B}\right)$. We then uniformly sample a source and target image pair $\left(I^{A},I^{B}\right)$ with $I^{A}\in P_{A},I^{B}\in P_{B}$ and a source and target image pair $\left(I^{A}_{L},I^{B}_{L}\right)$ with $I^{A}_{L}\in P_{A},I^{B}_{L}\in P_{B}$ for similarity loss computations. The pair $\left(I^{A},I^{B}\right)$ is used for transformation map prediction, and $\left(I_{L}^{A},I_{L}^{B}\right)$ is used for similarity loss calculation. The resulting training loss is
$\mathcal{L}=\mathcal{L}_{\operatorname{sim}}\left(I_{L}^{A}\circ\Phi_{\theta}\left[I^{A},I^{B}\right],I_{L}^{B}\right)$ $+\mathcal{L}_{\operatorname{sim}}\left(I_{L}^{B}\circ\Phi_{\theta}\left[I^{B},I^{A}\right],I_{L}^{A}\right)+$ (4)
$+\lambda\left\|\nabla\left(\Phi_{\theta}\left[I^{A},I^{B}\right]\circ\Phi_{\theta}\left[I^{B},I^{A}\right]\right)-\mathbf{I}\right\|_{F}^{2}.$
To explore the effects of choosing the modalities for image similarity calculations we experiment with both sampling $I_{L}^{A}$ and $I_{L}^{B}$ randomly or restricting the random sampling to picking the same modality for $I_{L}^{A}$ and $I_{L}^{B}$. Similar to our baseline approach, we use ($1-\text{LNCC}^{2}$) as our similarity measure. We train our model with the same hyperparameters as for uniGradICON. We set $\lambda=1.5$.
3.3 Dataset
We created a comprehensive training dataset by combining monomodal and multimodal datasets across anatomical regions; see Tab. 1 for details. We extend the uniGradICON [47] training dataset which contains only monomodal datasets. We also used additional modalities available in the uniGradICON datasets. Further, we added new datasets containing a wide range of brain MRI sequences (T1w, T1ce, T2w, FLAIR), contrasts derived from diffusion tensors (fractional anisotropy-FA, mean diffusivity-MD), CT-T1w abdomen MRIs, and DIXON MRIs for fat and water covering anatomical regions across the entire body from neck to knee. Our final corpus is composed of 16 different datasets, contains 5 different anatomical regions (lung, knee, brain, abdomen, pancreas) in addition to a whole body MR dataset, and 12 different image modalities (T1w, T1ce, T2w, T2, FLAIR, DESS, FA, MD, CT, CBCT, Fat/Water DIXON).
Table 1: Datasets used for training and testing.
Dataset Anatom. # of # of Type Modality Label % Training % Finetuning
region patients pairs Randomization Set Set
COPDGene [40] Lung 899 899 Intra-pat. CT 2.12 8.33
OAI [38] Knee 2532 7,398,400 Inter-pat. DESS/T2 MRI 6.38 12.5
HCP [50] Brain 1076 4,605,316 Inter-pat. T1w/T2w MRI 6.38 8.33
L2R-Abdomen [56] Abdomen 30 450 Inter-pat. CT 6.38 6.25
BratsReg [4] Brain 140 2,240 Intra-pat. T1w/T1ce/T2w/FLAIR MRI 21.27 8.33
ABCD [7] Brain 302 364,816 Inter-pat. FA/MD 6.38 0
L2R-AbdomenMRCT [12, 2, 32, 14] Abdomen 97 11,025 Inter-pat. CT/T1w MRI 12.76 6.25
UK Biobank [46] Neck-to-Knee 90 194,400 Inter-pat. Fat/Water DIXON 38.29 27.08
L2R-ThoraxCBCT-train [27, 28] Lung 14 1,764 Inter-pat. CT/CBCT 0 8.33
Pancreatic-CT-CBCT-SEG [24] Pancreas 40 720 Intra-pat. CT/CBCT 0 6.25
ABCD [7] Brain 307 1,483,524 Inter-pat. T1w/T2w/FA/MD 0 8.33
Dirlab-COPDGene [8] Lung 10 10 Intra-pat. CT - 0 0
OAI-test [38] Knee 301 301 Inter-pat. DESS MRI - 0 0
HCP-test [50] Brain 32 100 Inter-pat. T1w/T2w MRI - 0 0
L2R-NLST-val [1, 12] Lung 10 10 Intra-pat. CT - 0 0
L2R-OASIS-val [33, 25] Brain 20 19 Inter-pat. T1w MRI - 0 0
IXI-test [^2] https://brain-development.org/ixi-dataset/ Brain 115 115 Atlas-pat. T1w MRI - 0 0
L2R-ThoraxCBCT-val [27, 28] Lung 3 6 Intra-pat. CT/CBCT - 0 0
L2R-AbdomenMRCT-val [12, 2, 32, 14] Abdomen 2 3 Intra-pat. CT/T1w MRI - 0 0
UK Biobank-test [46] Neck-to-Knee 10 360 Inter-pat. Fat/Water DIXON - 0 0
Pancreatic-CT-CBCT-SEG [24] Pancreas 40 80 Intra-pat. CT/CBCT - 0 0
Data augmentation. We utilize affine data augmentation which randomly flips the input images and applies random affine transforms, as described in [48]. Further, inverting 0-1 normalized CT scans ($1-\text{CT}$) may enhance the segmentation accuracy of CT images when using a network trained on T1w MRIs [18]. This indicates that inverted CT scans may more strongly resemble T1w MRIs. Consequently, we integrate inverted CT scans into our training L2R-Abdomen and L2R-AbdomenMRCT datasets which already include CT images. During finetuning, we only apply random affine augmentation and do not use CT inversion to further fit our model on real image modalities.
Data balancing. The number of patients, scans, and provided modalities varies across datasets. To ensure a balanced dataset with respect to the number of possible modality-anatomical region combinations, we start by randomly selecting 4,000 image pairs from each dataset. These pairs are then assigned weights based on the dataset they belong to, ensuring an equal representation of observations from each (modality/region) combination during training. We then sample 4,000 3D image pairs per epoch using weighted sampling, consistent with the number of pairs per epoch used in uniGradICON [47]. For finetuning, we recompute our weights to account for the additional finetuning datasets; further, we use an anatomic-region-based weighting strategy that balances the number of seen anatomical regions by equally weighting anatomical regions for fine-tuning. Please refer to Tab. 1 for the diversity of the datasets and the percentages of each in the training and finetuning sets.
Data preprocessing. We clip the Hounsfield Units (HU) to the range [$-1000$, 1000] for all CT images and then normalize them to [0, 1]. For all MR images (T1w, T1ce, T2w, T2, FLAIR, DESS, DIXON, FA, MD), we clip the maximum intensity at the 99th percentile and then normalize them to [0, 1]. For the pancreatic CT-CBCT dataset, we follow the preprocessing steps outlined in [17]. We resize all images to a shape of [175, 175, 175] using trilinear interpolation. The spacing across the datasets varies; however, the input pairs (within a dataset) have the same spacing. During inference, we always evaluate our model on the original images by interpolating the transformation maps.
4 Results
We conduct experiments to assess multiGradICON’s performance and effectiveness compared to the existing monomodal foundational registration model, uniGradICON [47], and optimization-based registration method SyN [3], using default hyperparameters and MI as the similarity measure. We analyze monomodal and multimodal performance separately.
General hypothesis and questions. We hypothesize that multiGradICON can adapt to multimodal data while maintaining comparable monomodal performance to uniGradICON. However, our goal is for multiGradICON to be appropriate for multimodal and monomodal registration. Hence, for the monomodal setting, we seek answers to the questions: 1) How does the performance on monomodal datasets in the training set compare between our approach and uniGradICON?; 2) What is the performance difference in additional monomodal datasets that multiGradICON trained on compared to uniGradICON?; 3) Does multiGradICON generalize to unseen cases as well as uniGradICON?
Training design questions. We also investigate the impact of different factors such as training similarity loss selection ($1-\text{LNCC}$ or $1-\text{LNCC}^{2}$), instance optimization similarity loss selection ($1-\text{LNCC}^{2}$ or MIND-SSC), and training loss calculation strategy (baseline or label randomization). For this, we first train uniGradICON with $1-\text{LNCC}^{2}$ to investigate the effect of loss selection on generalization to multimodal pairs. Then, we introduce three variants of multiGradICON based on their image similarity loss calculation strategy: 1) the baseline multiGradICON-B approach which uses the same image pairing as input to the network and the loss; 2) multiGradICON-F which uses loss randomization but always samples from the same modality for the loss; 3) multiGradICON-R which also uses loss randomization but allows for sampling from different modalities. Finally, we obtain multiGradICON by further training our best-performing approach multiGradICON-R by including additional datasets to the training set (ThoraxCBCT, Pancreas, ACBD Diffusion (MD or FA)-Structure (T1w or T2w)). For this further training, we do not use $1-\text{CT}$ for data augmentation, we use lung-masked images for the COPDGene dataset, and we recompute the dataset weights (see Tab. 1). We report results for all our methods without instance optimization (w/o IO) or with 50 steps of IO using either $1-\text{LNCC}^{2}$ or MIND-SSC as similarity measures. Note that we perform all of the instance optimization operations on a given pair without any image similarity loss randomization, and we optimize the network parameters for the displacement field using an Adam optimizer with a learning rate of $2\times 10^{-5}$.
4.1 Performance on monomodal registration
Here, we discuss the performance of multiGradICON on monomodal datasets compared to uniGradICON. We split our evaluations into three categories based on the evaluation datasets: 1) Datasets that exist in both uni- and multiGradICON training sets; 2) Datasets that exist only in the multiGradICON training set; 3) Unseen datasets during training for uniGradICON and multiGradICON.
Table 2: Performance comparison on the monomodal datasets used for both uni- and multiGradICON training.
Lung Brain Abdomen Knee
COPDGene HCP Abdomen CTCT OAI
CT/CT (masked) CT/CT T1w/T1w CT/CT DESS/DESS
mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$
w/o IO SyN 8.20 0 15.18 0 75.8 0 25.2 0 65.7 0
uniGradICON 2.26 9.3e-5 6.71 5.7e-3 76.2 6.4e-5 48.3 3.1e-1 68.9 6.9e-2
uniGradICON-$\text{LNCC}^{2}$ 2.62 9.5e-5 6.59 1.3e-2 76.6 5.9e-5 49.8 3.2e-1 69.5 9.0e-2
multiGradICON - B 5.62 1.4e-3 6.34 3.3e-3 75.6 6.4e-5 39.2 2.2e-1 64.8 2.1e-2
multiGradICON - F 5.12 2.6e-4 5.56 2.0e-3 76.8 3.9e-5 40.5 5.1e-2 66.0 2.0e-2
multiGradICON - R 6.18 5.0e-4 6.60 3.6e-3 76.4 4.6e-5 40.2 2.7e-1 65.2 5.6e-2
multiGradICON 3.14 7.2e-4 5.85 4.0e-3 76.4 3.7e-5 39.5 8.6e-1 65.4 4.3e-2
1-LNCC2 uniGradICON 1.44 2.4e-4 2.80 1.3e-3 78.4 2.0e-4 52.9 9.4e-1 69.8 4.8e-2
uniGradICON-$\text{LNCC}^{2}$ 1.46 4.2e-4 2.97 1.7e-3 78.7 1.3e-4 53.4 8.9e-1 70.2 1.0e-2
multiGradICON - B 1.75 5.4e-5 2.65 3.6e-4 78.1 9.3e-5 46.5 9.7e-1 68.4 3.9e-2
multiGradICON - F 1.69 1.4e-4 2.48 5.3e-4 78.4 4.9e-5 48.1 6.2e-1 69.3 1.8e-2
multiGradICON - R 1.78 5.9e-5 2.91 3.8e-4 78.1 6.7e-5 47.6 7.2e-1 68.4 3.6e-2
multiGradICON 1.63 1.2e-4 2.92 5.2e-4 78.2 7.6e-5 46.9 6.5e-1 68.2 3.7e-2
MIND-SSC uniGradICON 1.77 2.6e-5 3.99 4.4e-5 77.6 3.7e-7 50.8 4.1e-1 69.3 4.9e-7
uniGradICON-$\text{LNCC}^{2}$ 1.80 6.7e-5 4.30 1.9e-4 77.7 1.6e-6 51.4 3.8e-1 69.7 9.8e-5
multiGradICON - B 2.22 0 3.79 7.4e-6 76.8 0 42.7 3.1e-1 66.6 0
multiGradICON - F 2.13 1.3e-5 3.39 5.5e-6 77.4 1.8e-7 44.5 8.6e-3 67.4 0
multiGradICON - R 2.25 0 3.92 7.4e-6 76.9 1.8e-7 44.4 3.6e-2 66.4 0
multiGradICON 2.03 1.3e-5 3.99 1.86e-6 77.1 7.4e-7 43.5 3.0e-2 66.4 0
Datasets used for both uni- and multiGradICON training. The uni- and multiGradICON training datasets both contain lung (COPDGene CT), brain (HCP T1w MRI), abdomen (L2R Abdomen CT), and knee (OAI MRI) images. Tab. 2 shows that uniGradICON outperforms multiGradICON-B,F,R on the lung, abdomen, and knee datasets based on the initial prediction without instance optimization. This performance difference is around $\sim$3 mm for COPDGene, $\sim$8.5Dice score for the abdomen, and $\sim$3Dice score for the knee dataset. This result is expected, as uniGradICON is exclusively trained on these datasets and thus has better expertise in these areas. Conversely, multiGradICON is trained on diverse datasets where these specific tasks have lower weight during training. This is further supported by the brain registration results, where multiGradICON-B,F,R show similar performance, with multiGradICON-F even outperforming uniGradICON-$\text{LNCC}^{2}$ by a $\sim$0.2Dice score. Since brain datasets are more prevalent in the training sets (e.g., ABCD and BratsReg), multiGradICON performs similarly to uniGradICON on brain registration. After finetuning with anatomical region-based sampling, we observe a performance improvement on the COPDGene dataset, which forms 2.1% of the training set but is sampled at 8.3% during finetuning. The performance on the remaining datasets remains similar since their percentages do not change drastically. We observe similar performance improvement on the unseen NLST lung dataset (see Sec. 4).
Instance optimization with $1-\text{LNCC}^{2}$ narrows the performance gap between uniGradICON and multiGradICON. The difference decreases to $\sim$0.19 mm for COPDGene, $\sim$6Dice score for the abdomen, and $\sim$1.6Dice score for the knee dataset. Instance optimization with MIND-SSC always under-performs instance optimization with $1-\text{LNCC}^{2}$ across all these datasets.
Different multimodal loss strategies perform differently for monomodal registration. Using the same modalities multiGradICON-F improves performance on the datasets to which this strategy is applied, compared to the baseline multiGradICON-B and random modality sampling in multiGradICON-R.
Lung masking also affects registration performance. We train our multiGradICON variants using full lung CT images, whereas uniGradICON uses masked images that are zeroed out outside the lung. We observe that a registration model trained without lung masking cannot generalize to register fine details of the lung even if we provide masked lungs during inference. Therefore, during the finetuning process, we further train our model with region of interest (ROI)-masked lung images. After that, we achieve approximately 3 mm improvement in mTRE on masked lung registration without IO. Additionally, we observe approximately 0.75 mm performance improvement on non-masked lung images, even though we do not incorporate any non-masked lung images in the finetuning process. We hypothesize that both the sampling amount in diverse datasets and ROI-masked training are crucial for achieving good registration performance.
Table 3: Comparison on monomodal datasets seen by multiGradICON but not by uniGradICON during training.
Brain Neck to Knee
HCP Brats-Reg UK Biobank
T2w/T2w T1w/T1w T2w/T2w T1ce/T1ce FLAIR/FLAIR WDIXON/WDIXON FDIXON/FDIXON
DICE(%) %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$
SyN 75.6 0 3.50 0 3.39 0 3.42 0 3.73 0 47.7 0 43.7 0
w/o IO uniGradICON 76.9 5.6e-4 3.27 1.0e-3 3.31 1.2e-3 3.24 1.4e-3 3.83 1.9e-3 42.2 8.1e-3 40.0 1.6e-2
uniGradICON-$\text{LNCC}^{2}$ 77.3 5.0e-4 3.22 0 3.21 0 3.13 0 3.79 0 42.4 1.6e-2 40.5 3.6e-2
multiGradICON - B 76.3 1.0e-4 3.10 6.1e-4 3.04 1.3e-3 2.91 6.9e-4 3.35 1.1e-3 43.6 3.9e-2 42.1 1.6e-2
multiGradICON - F 77.2 6.8e-5 2.94 7.4e-4 2.95 1.6e-3 2.73 9.2e-4 3.14 1.3e-3 45.5 1.8e-2 44.5 1.2e-2
multiGradICON - R 76.6 6.5e-5 3.07 1.2e-4 3.04 2.1e-3 2.87 1.3e-3 3.33 1.9e-3 43.6 4.4e-2 41.9 3.7e-2
multiGradICON 76.5 1.4e-4 3.06 8.8e-4 3.04 1.6e-3 2.90 1.2e-3 3.38 1.6e-3 43.8 2.3e-2 42.1 2.0e-2
1-LNCC2 uniGradICON 77.5 6.1e-4 2.93 7.7e-4 2.84 1.1e-3 2.48 9.4e-4 3.02 1.9e-3 47.0 8.5e-3 45.2 6.7e-3
uniGradICON-$\text{LNCC}^{2}$ 77.9 6.5e-4 2.92 8.8e-4 2.81 1.1e-3 2.45 9.1e-4 2.97 1.8e-3 46.8 1.3e-2 45.0 1.9e-2
multiGradICON - B 77.2 1.0e-4 2.94 7.3e-4 2.79 1.0e-3 2.50 7.7e-4 2.99 1.3e-3 47.9 5.7e-3 46.2 5.6e-3
multiGradICON - F 77.6 1.3e-4 2.92 8.1e-4 2.80 1.2e-3 2.49 9.9e-4 2.95 1.6e-3 48.6 5.2e-3 47.6 6.7e-3
multiGradICON - R 77.2 1.6e-4 2.92 8.8e-4 2.79 1.1e-3 2.47 9.7e-4 2.99 1.6e-3 47.1 6.0e-3 45.4 8.1e-3
multiGradICON 77.2 2.6e-4 2.92 9.7e-4 2.81 1.2e-3 2.48 1.0e-3 2.99 1.9e-3 47.3 6.9e-3 45.8 1.0e-2
MIND-SSC uniGradICON 77.2 7.8e-6 2.70 2.1e-6 2.54 0 2.20 0 2.62 0 45.0 0 43.1 1.0e-7
uniGradICON-$\text{LNCC}^{2}$ 77.5 2.0e-5 2.66 3.5e-5 2.48 0 2.14 0 2.55 2.6e-7 44.7 5.2e-8 43.1 0
multiGradICON - B 76.7 0 2.72 0 2.53 0 2.23 9.3e-7 2.62 0 45.6 0 44.1 6.7e-7
multiGradICON - F 77.2 7.4e-7 2.72 0 2.52 0 2.23 0 2.63 3.9e-7 46.8 0 46.0 0
multiGradICON - R 76.6 0 2.69 0 2.53 1.3e-7 2.23 1.3e-7 2.64 6.2e-6 44.8 2.1e-7 43.5 0
multiGradICON 76.6 7.4e-7 2.69 0 2.53 0 2.22 2.6e-7 2.62 1.3e-7 45.1 1.6e-7 43.7 0
Monomodal datasets that only exist in multiGradICON training. We additionally introduce new monomodal datasets to the multiGradICON training while retaining the existing uniGradICON training datasets. These new monomodal datasets comprise a wide variety of image modalities, such as T2w MRI, FLAIR, and DIXON, which have not been previously seen by uniGradICON. Tab. 3 shows the results across these datasets. Overall, the multiGradICON variants perform slightly better than uniGradICON on the Brats-Reg and UK Biobank datasets, since multiGradICON is trained on these domains. However, both approaches converge to similar performance after 50 steps of IO. These results demonstrate that even on previously unseen monomodal domains, uniGradICON remains a strong baseline, while multiGradICON performs better in initial predictions on the modalities it has seen during training.
Table 4: Performance comparison on unseen mono- and multimodal datasets for both uni- and multiGradICON.
Lung Brain Pancreas
NLST ThoraxCBCT IXI OASIS Pancreatic-CT-CBCT-SEG
CT/CT CT/CBCT T1w/T1w T1w/T1w CT/CBCT
mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$
w/o IO SyN 3.04 9.8-e1 57.4 0 64.5 1.0e-4 75.6 1.5e-2 78.2 0
uniGradICON 2.07 4.7e-4 57.0 4.7e-4 70.6 7.4e-3 79.0 8.9e-4 81.1 6.9e-2
uniGradICON-$\text{LNCC}^{2}$ 2.00 0 61.0 2.8e-3 69.7 2.1e-3 79.6 2.8e-3 81.0 8.1e-2
multiGradICON - B 2.74 0 58.1 3.6e-3 69.9 2.2e-4 78.6 1.7e-3 80.9 4.1e-2
multiGradICON - F 2.42 0 59.9 8.4e-3 71.6 2.9e-4 79.2 1.7e-3 80.5 2.3e-2
multiGradICON - R 2.66 0 58.9 3.8e-2 70.7 3.4e-4 78.5 2.3e-3 80.6 5.6e-2
multiGradICON 2.27 0 58.7 3.2e-3 71.0 1.8e-3 78.7 2.0e-3 81.8 2.1e-2
1-LNCC2 uniGradICON 1.77 8.7e-5 60.9 2.3e-1 70.4 1.5e-3 79.7 6.5e-3 82.2 2.4e-2
uniGradICON-$\text{LNCC}^{2}$ 1.76 4.8e-5 62.1 2.5e-2 70.8 1.6e-3 80.1 9.1e-3 82.0 4.2e-2
multiGradICON - B 1.84 3.1e-4 60.1 2.0e-2 70.8 1.0e-3 79.5 5.6e-3 82.0 1.9e-2
multiGradICON - F 1.82 2.1e-4 61.4 4.3e-1 70.7 1.1e-3 79.9 6.9e-3 82.3 7.9e-3
multiGradICON - R 1.86 2.3e-4 60.2 2.7e-1 70.7 1.4e-3 79.6 7.1e-3 82.1 8.8e-3
multiGradICON 1.79 2.7e-5 60.6 2.8e-1 71.0 1.8e-3 79.6 6.3e-3 82.2 8.3e-3
MIND-SSC uniGradICON 1.87 0 57.9 2.3e-2 71.7 1.6e-6 78.9 3.4e-5 82.0 1.1e-6
uniGradICON-$\text{LNCC}^{2}$ 1.84 0 58.3 1.8e-2 72.2 8.9e-7 79.3 0 81.8 1.5e-5
multiGradICON - B 1.99 0 59.4 9.1e-1 72.0 2.5e-7 78.6 3.1e-6 81.5 3.4e-6
multiGradICON - F 1.95 0 63.5 1.4e-3 71.7 1.0e-6 79.1 0 81.8 0
multiGradICON - R 2.03 0 63.0 6.6e-6 71.6 3.2e-6 78.7 2.3e-6 81.7 2.3e-7
multiGradICON 1.94 0 63.4 0 71.8 4.3e-6 78.2 0 81.9 0
Unseen monomodal datasets. We also test on datasets that are never seen during training. Tab. 4 shows performance metrics for lung CT-CT NLST and brain T1w MRI registrations from the IXI and OASIS datasets. For NLST, uniGradICON achieves an mTRE of 2.07 mm, whereas the multiGradICON-B,F,R variants show a range between 2.42 and 2.74 mm. We observe a performance improvement on the NLST dataset of approximately 0.4 mm after finetuning.
Instance optimization with $1-\text{LNCC}^{2}$ decreases the performance gap to $\sim$0.02 mm on the NLST dataset. However, for brain registration, multiGradICON shows similar performance to uniGradICON, outperforming it by a $\sim$0.6Dice score on the IXI dataset and underperforming it by a $\sim$0.1Dice score on the OASIS dataset. These results show that multiGradICON scales well to monomodal tasks, providing performance close to that of uniGradICON, which specializes in monomodal registration.
4.2 Performance on multimodal registration
In this section, we evaluate the multimodal registration performance of multiGradICON on several datasets including a wide variety of anatomical structures and modalities. We again investigate its performance on both seen and unseen datasets. We use uniGradICON as our main comparison model.
Multimodal training datasets. Our training dataset consists of several anatomical regions and modalities such as T1w, T1ce, T2w, FLAIR brain MRIs, CT abdominal scans, and fat and water-weighted DIXON images (see Tab. 1 for details). Tab.  5 shows registration performances for the HCP, Brats-Reg, Abdomen MR/CT, and UK Biobank datasets. In both the HCP and Brats-Reg datasets, we observe that uniGradICON fails to register pairs that contain images with large appearance differences. For instance, uniGradICON cannot register pairs containing T2w images, even with instance optimization. This is one of the key problems that we aim to solve with multiGradICON. We observe a significant performance improvement on the HCP ($\sim$59Dice score improvement), Brats-Reg, and MR-CT ($\sim$11Dice score improvement) datasets without instance optimization with our multiGradICON approach.
Table 5: Performance comparison on the multimodal datasets used for multiGradICON training.
Brain Abdomen Neck to Knee
HCP Brats-Reg Abdomen MRCT UK Biobank
T1w/T2w T1w/T2w T1w/T1ce T1w/FLAIR T2w/T1ce T2w/FLAIR T1ce/FLAIR MR/CT FDIXON/WDIXON
DICE(%) %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ mTRE %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$ DICE(%) %$\%|J|_{<0}$
w/o IO SyN 71.9 0 3.74 0 3.62 0 3.91 0 3.74 0 3.96 0 3.88 0 45.0 0 33.9 0
uniGradICON 10.0 2.6e-3 13.11 3.7e-3 4.16 2.1e-3 6.38 2.3e-3 12.30 4.6e-3 8.12 5.0e-3 6.92 3.9e-3 50.0 4.0e-2 17.5 2.4e-2
uniGradICON-$\text{LNCC}^{2}$ 14.5 8.6e-4 13.10 4.9e-3 4.08 4.2e-3 6.33 3.2e-3 12.24 6.8e-3 7.95 6.0e-3 6.83 6.3e-3 49.7 2.6e-2 18.1 4.4e-2
multiGradICON - B 69.2 8.0e-4 3.60 1.7e-3 3.84 8.1e-4 4.60 1.2e-3 3.73 2.4e-3 4.95 2.9e-3 4.95 2.3e-3 58.1 3.4e-2 19.7 1.2e-2
multiGradICON - F 40.1 1.0e-3 9.10 2.2e-3 3.66 1.2e-3 5.06 1.4e-3 8.56 2.8e-3 6.50 3.2e-3 5.60 2.3e-3 58.2 6.0e-2 20.1 1.2e-2
multiGradICON - R 68.2 9.0e-4 3.58 2.3e-3 3.78 1.5e-3 4.48 2.1e-3 3.70 2.9e-3 4.90 3.8e-3 4.76 3.1e-3 61.1 8.4e-2 19.8 2.9e-2
multiGradICON 69.0 7.9e-4 3.64 2.5e-3 3.83 1.6e-3 4.68 1.9e-3 3.78 3.4e-3 5.05 4.1e-3 5.04 3.3e-3 61.8 4.0e-2 19.7 1.3e-2
1-LNCC2 uniGradICON 7.8 3.3e-3 12.06 3.2e-3 3.63 1.6e-3 5.13 2.0e-3 11.69 3.8e-3 6.81 4.8e-3 5.71 3.3e-3 75.1 6.9e-1 17.0 6.9e-3
uniGradICON-$\text{LNCC}^{2}$ 8.3 2.4e-3 11.94 2.7e-3 3.59 1.8e-3 5.05 1.9e-3 11.48 3.2e-3 6.60 4.5e-3 5.63 3.2e-3 80.3 5.8e-1 17.4 8.6e-3
multiGradICON - B 72.8 6.3e-4 3.27 7.4e-4 3.56 1.3e-3 4.28 1.1e-3 3.34 1.1e-3 4.58 2.5e-3 4.76 2.1e-3 73.3 4.4e-1 18.7 5.5e-3
multiGradICON - F 71.9 4.9e-4 3.34 9.5e-4 3.55 1.5e-3 4.29 1.4e-3 3.39 1.5e-3 4.57 3.1e-3 4.78 2.6e-3 74.8 3.0e-1 18.9 6.9e-3
multiGradICON - R 72.8 5.7e-4 3.30 9.4e-4 3.55 1.5e-3 4.24 1.4e-3 3.37 1.4e-3 4.55 2.8e-3 4.70 2.5e-3 74.8 1.6e-1 18.6 6.6e-3
multiGradICON 73.1 9.3e-4 3.32 1.1e-3 3.56 1.8e-3 4.27 1.6e-3 3.38 1.8e-3 4.60 3.2e-3 4.72 3.0e-3 73.3 2.0e-1 18.6 8.0e-3
MIND-SSC uniGradICON 39.8 1.4e-5 2.66 0 2.61 8.7e-7 2.81 1.9e-6 2.55 6.7e-8 2.88 6.7e-8 2.75 7.3e-7 75.4 1.9e-4 20.2 0
uniGradICON-$\text{LNCC}^{2}$ 53.0 9.5e-6 2.58 6.7e-8 2.55 6.7e-8 2.74 2.7e-7 2.47 6.7e-8 2.80 0 2.68 6.7e-8 77.4 1.7e-3 20.8 3.1e-7
multiGradICON - B 74.5 0 2.70 0 2.61 8.0e-7 2.83 6.7e-8 2.62 0 2.94 0 2.79 3.3e-7 68.1 5.0e-3 23.3 0
multiGradICON - F 73.0 9.3e-7 2.72 0 2.62 0 2.87 6.7e-8 2.65 2.7e-7 2.94 2.0e-7 2.82 0 70.9 3.2e-4 25.2 0
multiGradICON - R 74.3 0 2.73 0 2.62 6.7e-8 2.88 6.7e-7 2.66 0 2.97 0 2.83 6.7e-8 70.3 1.0e-3 23.7 0
multiGradICON 74.4 1.4e-6 2.70 1.3e-7 2.62 1.2e-6 2.85 1.6e-6 2.62 0 2.94 1.3e-7 2.80 4.7e-7 70.8 2.7e-4 23.3 0
On the other hand, the UK Biobank fat-water weighted DIXON registration is challenging due to the different underlying information across modalities. Although this is a clinically irrelevant scenario since they are acquired together, this result demonstrates the limitation of our approach. We will address this type of registration, where the pairs share the same anatomies but capture entirely different properties, in future work by incorporating semantic information. Additionally, we note that we obtain UK Biobank segmentations using MRSegmentator [19] and evaluate our models directly based on MRSegmentator predictions. Therefore, we can only provide a silver-standard performance metric that may also include possible segmentation errors arising from the MRSegmentator.
We observe that multiGradICON-F cannot generalize to multimodal pairs compared to the other multiGradICON variants. It shows poor performance on several registration tasks such as T1w-T2w, T1ce-T2w, and T1ce-FLAIR brain registration. Although multiGradICON-F performs better than uniGradICON, these results suggest that sampling the same modality for similarity loss calculation causes a performance drop on multimodal datasets, while it improves the performance on monomodal pairs, as we discussed in Sec. 4.1. Additionally, we do not observe any significant multimodal performance improvement in uniGradICON when $1-\text{LNCC}^{2}$ is used as similarity loss during training. This indicates that using appropriate losses for multimodal registration may not lead to multimodal generalization if multimodal datasets were not used during training. We hypothesize that the diversity of the training dataset is more important for generalization.
Moreover, we observe improved registration results with instance optimization (IO) using the MIND-SSC loss in the Brats-Reg dataset compared to IO with $1-\text{LNCC}^{2}$. The Brats-Reg dataset consists of pre-operative and follow-up brain scans that include tumors with varying shapes and resections where each modality captures different tumor properties. As MIND-SSC provides better alignment for these inconsistent structures than $1-\text{LNCC}^{2}$, we conclude that task-specific similarity loss selection can be important for good multimodal registration.
Unseen multimodal datasets. During the training of multiGradICON, we initially did not include the ThoraxCBCT and pancreatic CT-CBCT datasets in our training set. Although the performances of uniGradICON and multiGradICON-B,F,R on these datasets are close to each other, uniGradICON underperforms by $\sim$2Dice scores on the ThoraxCBCT dataset and outperforms by $\sim$0.5Dice score on the pancreatic CT-CBCT dataset. However, finetuning that incorporates the pancreatic dataset leads to a better Dice score for multiGradICON which outperforms uniGradICON by $\sim$0.7Dice score. The performance on the ThoraxCBCT dataset remains similar even when it is included in the training set during finetuning. We hypothesize that our model already converges on CT images during the initial training and that additional CBCT images do not affect the performance since they look sufficiently similar to CT images.
5 Conclusion
We developed multiGradICON, a universal deep network for mono- and multimodal registration. multiGradICON extends uniGradICON, the first deep medical registration network for monomodal registration across different anatomies. We observed that uniGradICON remains a strong baseline for monomodal registration but multiGradICON shows improved performance for multimodal registration, in particular, when modalities look so different that not even instance optimization recovers good registrations for uniGradICON. We also demonstrated that similarity loss randomization can bring multimodal registration benefits. Although multiGradICON showed encouraging performance, there are many different avenues for future improvements. For example, we only investigated using $1-\text{LNCC}^{2}$ as a training loss, while many other multimodal similarity measures exist. Further, multiGradICON cannot reliably register between DIXON fat and water images. Although this is somewhat expected, it points to the existence of multimodal datasets that share underlying anatomy but have such large appearance differences that a similarity measure such as $1-\text{LNCC}^{2}$ is insufficient. Using segmentations as part of the image similarity loss may show benefits in such cases. In general, training with segmentations [55] or simulated data for the network input and the similarity loss could be a fruitful avenue for future work. Lastly, just like uniGradICON, multiGradICON is built on top of the exact same network architecture as GradICON. Exploring networks with increased capacity and further increasing training dataset sizes would be desirable.
6 Acknowledgements
[…]
Appendix 0.A Additional visualizations
Fig. 2 shows additional registration results for mulitGradICON for monomodal and multimodal registration without instance optimization. We observe that multiGradICON can register a wide variety of anatomies and image modalities; with some of the modality/anatomy pairings (e.g., for CT/MRI in the abdomen) being highly challenging.
Figure 2: Visualizations of monomodal and multimodal registration results for multiGradICON without instance optimization. The results demonstrate that multiGradICON can handle a wide variety of modalities and anatomies with smooth displacement fields. Note that these images are visualized in their interpolated shapes as they are provided to the network. DIXON images reproduced by kind permission of the UK Biobank®.